The sum of three consecutive even integers is 480. Find the three integers
Answer by JulietG(1812)

A + B + C = 480
B = A + 2
C = A + 4
.
Substitute the values
A + B + C = 480
A + (A+2) + (A+4) = 480
Add
3A + 6 = 480
Subtract 6 from each side
3A = 474
Divide each side by 3
A = 158
therefore B = 160 and C = 162
.
add them together to prove
158 + 160 + 162 does indeed equal 480
Success!